# Diversity-Aware Policy Optimization for Large Language Model Reasoning

### v1

@@ Abstract
The reasoning capabilities of large language models (LLMs) have advanced rapidly, particularly following the release of DeepSeek-R1, which has inspired a surge of research into data quality and reinforcement learning (RL) algorithms. Despite the pivotal role diversity plays in RL, its influence on LLM reasoning remains largely underexplored. To bridge this gap, this work presents a systematic investigation into the impact of diversity in RL-based training for LLM reasoning, and proposes a novel diversity-aware policy optimization method. Across evaluations on 12 LLMs, we observe a strong positive correlation between the solution diversity and Potential@k (a novel metric quantifying an LLM’s reasoning potential) in high-performing models. This finding motivates our method to explicitly promote diversity during RL training. Specifically, we design a token-level diversity and reformulate it into a practical objective, then we selectively apply it to positive samples. Integrated into the R1-zero training framework, our method achieves a 3.5% average improvement across four mathematical reasoning benchmarks, while generating more diverse and robust solutions.
@@ 1 Introduction
Recently, the reasoning capabilities of large language models (LLMs) have made remarkable progress, with significant improvements showcased by OpenAI-o1 [31], DeepSeek-R1 [12], and Kimi-k1.5 [39]. Among these advancements, two key innovations have contributed significantly: First, the adoption of a rule-based reward system significantly streamlines the training process by focusing exclusively on rewarding correct final answers and proper output formats, thereby eliminating the complexity associated with process-based reward models [26, 43]. Second, the introduction of a lightweight reinforcement learning (RL) algorithm [12, 39] removes the need for a separate critic model, substantially reducing computational overhead and accelerating the training process. The success of DeepSeek-R1 has attracted numerous follow-up studies [55], which broadly fall into two categories. The first category focuses on improving the quality of training data [29, 14, 19, 2], emphasizing rigorous data set curation through filtering, deduplication, and verification. The second category refines RL algorithms, including detailed optimizations for PPO-based methods such as VCPPO [49] and VAPO [48]; enhancements to GRPO for stability and speed, such as DAPO [47], Dr.GRPO [27] and SRPO [56]; as well as alternative approaches such as REINFORCE++ [18].
While RL has been extensively applied to LLM reasoning, the role of diversity remains largely unexplored in this context, even though it plays a crucial role in RL research[17, 10, 32, 7, 33, 28, 57, 11, 51, 58, 5]. In traditional RL tasks, incorporating diversity is widely recognized to facilitate exploration by promoting the selection of more stochastic policies, which helps the policy escape local optima and accelerate the convergence of training. This hypothesis has been experimentally validated in previous work [17, 10, 32]. Beyond empirical evidence, theoretical analyses suggest that policies with higher entropy (a measure of diversity) can smooth the optimization landscape [1]. These findings naturally lead us to ask the following question: Is promoting diversity essential during RL training for LLM reasoning?
@@
The empirical findings motivate us to promote diversity during RL training for LLM reasoning. A commonly used approach for this goal is entropy regularization. However, directly increasing the average entropy of LLM outputs can introduce length bias, as longer responses inherently exhibit higher entropy. To address this, we introduce a token-level diversity metric and reformulate the diversity objective into a practical form. Moreover, promoting diversity often entails a quality-diversity trade-off. To mitigate this, we strategically apply diversity enhancement only to positive samples, thereby enriching solution diversity while preserving training stability. This design is akin to fostering diversity in high-quality policies in population-based RL training, ensuring that exploration is guided by task-relevant performance criteria [44]. Finally, we integrate our diversity objective into the R1-zero training method and evaluate the enhanced approach across $4$ mathematical reasoning benchmarks. Experimental results demonstrate a 3.5% average performance gain over standard R1-zero training, while our method can generate more diverse solutions.
@@ 2.1 RL for LLMs
Formally, in the context of LLM generation for mathmatical problem-solving, where each prompt is a question, we define the prompt as $q\in\mathcal{Q}$, where $\mathcal{Q}$ represents the set of all possible questions. The set of all potential text outputs $o$ forms an action space $\mathcal{O}$. Each output $o$ consists of tokens, denoted as $o:=(o^{1},o^{2},...,o^{t},...)$. To generate an output, a policy $\pi_{\theta}(\cdot|q)$ parameterized by $\theta$ is employed, which generates the output according to the distribution:
@@ 2.2 Reinforcement Learning algorithm
The R1-zero training method proposed by DeepSeek-R1 [12] has attracted significant research attention due to its computational efficiency and effectiveness. In our work, we adopt this training method as our backbone. R1-zero incorporates two key innovations: the GRPO algorithm [36] and a rule-based reward function. In this section, we introduce both components.
@@
[…]
@@ 3 Correlation between LLMs’ reasoning potential and solution diversity
The role of diversity has long been established as critical in traditional RL tasks. Numerous studies [17, 10, 32, 7, 33] have shown that promoting diversity can enhance the final quality of the policy. However, its impact in the realm of RL for LLM reasoning still remains under-explored. In this section, we investigate the relationship between solution diversity and the reasoning abilities of LLMs on mathematical benchmarks. We adopt the equation diversity in prior work [45] to quantify the variety of solutions generated for mathematical problem-solving. For reasoning ability, we introduce a novel metric to evaluate an LLM’s training potential (related to the performance gain achieved after RL training).
@@
We evaluate 12 LLMs on the MATH benchmark [15]. For each question, we calculate: (1) Pass@1 accuracy using greedy decoding, and (2) Diversity with (3) Potential@k, both evaluated from $16$ sampled responses (temperature=0.9).
For diversity, we adopt the metric (denoted as Div-Equ) from prior work [45], which measures the ratio of distinct equations among the responses:
[…]
@@
The results are shown in Figure 1(a). The results show a bifurcated pattern: For LLMs with limited reasoning ability (Pass@1 $<0.4$), we observe no significant relationship between solution diversity and model potential. For stronger performers (Pass@1 $>0.4$), a clear positive correlation emerges between these metrics. Linear regression on this high-performing subset yields $R^{2}=0.81$, confirming a strong predictive relationship where increased diversity corresponds to higher model potential.
@@ 4 Diversity-aware policy optimization
Building on the insights from Section 3, in this section, we introduce an entropy-based diversity and propose its targeted application to positive samples during policy optimization for LLM reasoning. We incorporate this diversity objective into the R1-zero training method [12], which employs the GRPO algorithm with the reward function defined in Section 2.2. We refer to this enhanced approach as R1-zero-Div.
@@ 4.2 Promoting diversity on positive samples
This is akin to fostering diversity in high-quality policies in population-based RL training [44], while we focus on positive samples rather than policies here. Beyond intuitive justification, we further justify this design by analyzing the gradient on each token.
@@
$-\nabla_{{\theta}}\pi_{\theta}(o^{t}|q,o^{<t})\log\pi_{\theta}(o^{t}|q,o^{<t})=-[1+\log\pi_{\theta}(o^{t}|q,o^{<t})]\cdot\nabla_{{\theta}}\pi_{\theta}(o^{t}|q,o^{<t}).$ (11)
Hence, for tokens with small probabilities (in that case $\pi_{\theta}(o^{t}|q,o^{<t})<e^{-1}$, and this holds for most tokens since the sum of probability is equal to $1$), the gradient aligns with $\nabla_{{\theta}}\pi_{\theta}(o^{t}|q,o^{<t})$. This suggests that the diversity component’s gradient actively promotes increasing the probability of low-probability tokens, which inherently offer substantial growth potential. However, this tendency is undesirable for negative samples. Thus, excluding diversity enhancement for negative samples mitigates conflicts between solution quality and diversity. A visual illustration is provided in Figure 1(b). Moreover, the experimental results in Section 5.3 and Appendix E.1 further support our design.
@@ 5.1 Experimental setup
We choose Qwen2.5-Math-7B [46] as our base model, which is commonly used for mathematical reasoning benchmarks [52, 59, 20]. Additionally, we conduct an ablation study using Qwen2.5-Math-1.5B [46] to assess the effectiveness of our approach in smaller LLMs.
@@
We selected $4$ mathematical benchmarks to evaluate the models’ reasoning abilities: GSM8K [6], MATH500 [15], Olympiad Bench [13], and College Math [38]. Each contains at least 500 data points for testing. We excluded some commonly used mathematical benchmarks that provide limited data, e,g, AIME24  [^1] https://huggingface.co/datasets/Maxwell-Jia/AIME_2024 with 30 items, as they can lead to unstable and biased evaluation outcomes. We train the base model on the GSM8K training set and then evaluate on the $4$ benchmarks.
@@
The most pertinent baselines for comparison are the base model itself and the base model trained via R1-zero. Additionally, we incorporate the latest prominent "R1-zero-Like" models with similar backbones for reference: SimpleRL-Zoo [52], PRIME-Zero-7B [8]. It is important to note that these methods are trained with different computational resources and datasets, making direct comparisons challenging. Our approach is designed to enhance diversity rather than compete directly with these methods. In fact, our method is compatible with and can be integrated into these existing approaches.
@@
For R1-zero-Div, we train the base model on the GSM8K training set using the loss function in Equation 4.2, with a learning rate of $3\times 10^{-6}$ and the AdamW optimizer. During rollout, we sample $6$ responses with a temperature of $0.9$ and train for $2$ epochs. Our implementation is built on TRL [42] and runs on 8$\times$×A6000 GPUs. For R1-zero, we maintain identical settings to R1-zero-Div but exclude the diversity objective. For other baselines, we evaluate open-sourced models downloaded from Hugging Face [^2] https://huggingface.co, following the settings recommended in their original papers. Additional implementation details are provided in Appendix D.
@@ 5.2 Main results
We evaluate the reasoning performance using Pass@1 accuracy, as shown in Table 1. We also report the performance against training steps in Figure 2. In our experiment, R1-zero-Div demonstrates superior performance compared to R1-zero, achieving an average improvement of $3.5\%$. Despite being trained with limited computational resources (discussed in Appendx B), R1-zero-Div achieves comparable results to state-of-the-art methods (SimpleRL-Zoo and Eurus-2-7B-PRIME). These results suggest that promoting diversity on positive samples in training can effectively enhance the model’s reasoning capabilities.
@@
We empirically demonstrate that R1-zero-Div produces more diverse solutions than other RL-finetuning baselines. Our evaluation on the GSM8K test set generates 5 responses for each of 1,319 questions, measuring diversity through three metrics: Div-Equ, and two additional metrics in prior work [22]: (1) N-gram diversity (proportion of distinct n-grams per response, capturing intra-diversity) and (2) Self-BLEU diversity (100 minus Self-BLEU score, capturing inter-diversity). All metrics range from 0 to 100, with higher values indicating greater diversity. As shown in Table 2, while RL fine-tuning methods significantly reduce diversity (compared to the base model), R1-zero-Div effectively preserves diversity. We further provide concrete examples in Appendix E showing that R1-zero-Div generates distinct solutions for the same question.
@@
Method GSM8K MATH500 Olympiad Bench College Math Avg
Base 57.6 51.8 16.1 21.4 36.7
Base-R1-zero 88.7 75.6 38.2 46.2 62.2
Base-R1-zero-Div (Ours) 91.7 78.2 40.1 47.6 64.4
Base-SimpleRL-Zoo 90.2 80.0 39.0 47.2 64.1
Base-Eurus-2-7B-PRIME 88.4 78.2 39.4 46.6 63.2
Figure 2: Pass@1 Accuracy (on test set) against the training steps.
Table 2: Diversity of different methods on GSM8K test set
@@
Base 92.26 29.29 85.98
Base-Eurus-2-7B-PRIME 60.86 24.08 48.20
Base-SimpleRL-Zoo 74.89 25.41 49.32
Base-R1-zero 75.02 27.75 56.00
Base-R1-zero-Div (Ours) 79.29 29.60 58.89

### v2

@@ Abstract
The reasoning capabilities of large language models (LLMs) have advanced rapidly, particularly following the release of DeepSeek-R1, which has inspired a surge of research into data quality and reinforcement learning (RL) algorithms. Despite the pivotal role diversity plays in RL, its influence on LLM reasoning remains largely underexplored. To bridge this gap, this work presents a systematic investigation into the impact of diversity in RL-based training for LLM reasoning, and proposes a novel diversity-aware policy optimization method. Across evaluations on 12 LLMs, we observe a strong positive correlation between the solution diversity and Potential@k (a novel metric quantifying an LLM’s reasoning potential) in high-performing models. This finding motivates our method to explicitly promote diversity during RL training. Specifically, we design a token-level diversity and reformulate it into a practical objective, then we selectively apply it to positive samples. Integrated into the R1-zero training framework, our method achieves a 3.5% average improvement across four mathematical reasoning benchmarks, while generating more diverse and robust solutions. The code is available at https://github.com/nigelyaoj/R1_zero_Div.
@@ 1 Introduction
Recently, the reasoning capabilities of large language models (LLMs) have made remarkable progress, with significant improvements showcased by OpenAI-o1 (openai-o1, ), DeepSeek-R1 (guo2025deepseek, ), and Kimi-k1.5 (team2025kimi, ). Among these advancements, two key innovations have contributed significantly: First, the adoption of a rule-based reward system significantly streamlines the training process by focusing exclusively on rewarding correct final answers and proper output formats, thereby eliminating the complexity associated with process-based reward models (lightman2023let, ; wang2023math, ). Second, the introduction of a lightweight reinforcement learning (RL) algorithm (guo2025deepseek, ; team2025kimi, ) removes the need for a separate critic model, substantially reducing computational overhead and accelerating the training process. The success of DeepSeek-R1 has attracted numerous follow-up studies (zhang2025100, ), which broadly fall into two categories. The first category focuses on improving the quality of training data (meng2023deepscaler, ; he2025deepmath, ; hu2025openreasonerzeroopensourceapproach, ; albalak2025big, ), emphasizing rigorous data set curation through filtering, deduplication, and verification. The second category refines RL algorithms, including detailed optimizations for PPO-based methods such as VCPPO (yuan2025s, ) and VAPO (yuan2025vapo, ); enhancements to GRPO for stability and speed, such as DAPO (yu2503dapo, ), Dr.GRPO (liu2025understanding, ) and SRPO (zhang2025srpo, ); as well as alternative approaches such as REINFORCE++ (hu2025reinforce++, ).
While RL has been extensively applied to LLM reasoning, the role of diversity remains largely unexplored in this context, even though it plays a crucial role in RL research(hong2018diversity, ; eysenbach2018diversity, ; parker2020effective, ; conti2018improving, ; peng2020non, ; masood2019diversity, ; zhang2019learning, ; ghasemi2021multiple, ; zahavy2021discovering, ; zhou2022continuously, ; cideron2020qd, ; yangdiverse, ). In traditional RL tasks, incorporating diversity is widely recognized to facilitate exploration by promoting the selection of more stochastic policies, which helps the policy escape local optima and accelerate the convergence of training. This hypothesis has been experimentally validated in previous work (hong2018diversity, ; eysenbach2018diversity, ; parker2020effective, ). Beyond empirical evidence, theoretical analyses suggest that policies with higher entropy (a measure of diversity) can smooth the optimization landscape (Ahmed_Roux_Norouzi_Schuurmans_2019, ). These findings naturally lead us to ask the following question: Is promoting diversity essential during RL training for LLM reasoning?
@@
The empirical findings motivate us to promote diversity during RL training for LLM reasoning. A commonly used approach for this goal is entropy regularization. However, directly increasing the average entropy of LLM outputs can introduce length bias, as longer responses inherently exhibit higher entropy. To address this, we introduce a token-level diversity metric and reformulate the diversity objective into a practical form. Moreover, promoting diversity often entails a quality-diversity trade-off. To mitigate this, we strategically apply diversity enhancement only to positive samples, thereby enriching solution diversity while preserving training stability. This design is akin to fostering diversity in high-quality policies in population-based RL training, ensuring that exploration is guided by task-relevant performance criteria (wu2023quality, ). Finally, we integrate our diversity objective into the R1-zero training method and evaluate the enhanced approach across $4$ mathematical reasoning benchmarks. Experimental results demonstrate a 3.5% average performance gain over standard R1-zero training, while our method can generate more diverse solutions.
@@ 2.1 RL for LLMs
[…]
@@ 2.2 Reinforcement Learning algorithm
The R1-zero training method proposed by DeepSeek-R1 (guo2025deepseek, ) has attracted significant research attention due to its computational efficiency and effectiveness. In our work, we adopt this training method as our backbone. R1-zero incorporates two key innovations: the GRPO algorithm (shao2024deepseekmath, ) and a rule-based reward function. In this section, we introduce both components.
@@
In line with DeepSeek-R1 (guo2025deepseek, ), we implement two types of rule-based rewards: accuracy rewards and format rewards. The accuracy reward model assesses whether the response is correct by comparing the predicted answer to the golden reference answer, while the format reward model ensures that the final answer is presented in a $\backslash{\rm boxed}\{\}$ format for reliable verification.
@@ 3 Correlation between LLMs’ reasoning potential and solution diversity
The role of diversity has long been established as critical in traditional RL tasks. Numerous studies (hong2018diversity, ; eysenbach2018diversity, ; parker2020effective, ; conti2018improving, ; peng2020non, ) have shown that promoting diversity can enhance the final quality of the policy. However, its impact in the realm of RL for LLM reasoning still remains under-explored. In this section, we investigate the relationship between solution diversity and the reasoning abilities of LLMs on mathematical benchmarks. We adopt the equation diversity in prior work (wu2024progress, ) to quantify the variety of solutions generated for mathematical problem-solving. For reasoning ability, we introduce a novel metric to evaluate an LLM’s training potential (related to the performance gain achieved after RL training).
@@
We evaluate 12 LLMs on the MATH benchmark (hendrycks2021measuring, ). For each question, we calculate: (1) Pass@1 accuracy using greedy decoding, and (2) Diversity with (3) Potential@k, both evaluated from $16$ sampled responses (temperature=0.9).
For diversity, we adopt the metric (denoted as Div-Equ) from prior work (wu2024progress, ), which measures the ratio of distinct equations among the responses:
$\text{Div-Equ}:=\frac{1}{N}\sum_{i=1}^{N}\frac{|\mathcal{U}_{i}|}{|\mathcal{A}_{i}|},$ (5)
@@
The results are shown in Figure 1(a). The results show a bifurcated pattern: For LLMs with limited reasoning ability (Pass@1 <0.4<0.4), we observe no significant relationship between solution diversity and model potential. For stronger performers (Pass@1 >0.4>0.4), a clear positive correlation emerges between these metrics. Linear regression on this high-performing subset yields $R^{2}=0.81$, confirming a strong predictive relationship where increased diversity corresponds to higher model potential.
@@ 4 Diversity-aware policy optimization
Building on the insights from Section 3, in this section, we introduce an entropy-based diversity and propose its targeted application to positive samples during policy optimization for LLM reasoning. We incorporate this diversity objective into the R1-zero training method (guo2025deepseek, ), which employs the GRPO algorithm with the reward function defined in Section 2.2. We refer to this enhanced approach as R1-zero-Div.
@@ 4.2 Promoting diversity on positive samples
This is akin to fostering diversity in high-quality policies in population-based RL training (wu2023quality, ), while we focus on positive samples rather than policies here. Beyond intuitive justification, we further justify this design by analyzing the gradient on each token.
@@
$\nabla_{{\theta}}\pi_{\theta}(o^{t}|q,o^{<t})\log\pi_{\theta}(o^{t}|q,o^{<t})=-[1+\log\pi_{\theta}(o^{t}|q,o^{<t})]\cdot\nabla_{{\theta}}\pi_{\theta}(o^{t}|q,o^{<t}).$ (11)
Hence, for tokens with small probabilities (in that case $\pi_{\theta}(o^{t}|q,o^{<t})<e^{-1}$, and this holds for most of tokens since the sum of probability is equal to $1$), the gradient aligns with $\nabla_{{\theta}}\pi_{\theta}(o^{t}|q,o^{<t})$. This suggests that the diversity component’s gradient actively promotes increasing the probability of low-probability tokens, which inherently offer substantial growth potential. However, this tendency is undesirable for negative samples. Thus, excluding diversity enhancement for negative samples mitigates conflicts between solution quality and diversity. A visual illustration is provided in Figure 1(b). Moreover, the experimental results in Section 5.3 and Appendix E.1 further support our design.
@@ 5.1 Experimental setup
We choose Qwen2.5-Math-7B (Qwen7B) (yang2024qwen25mathtechnicalreportmathematical, ) as our base model, which is commonly used for mathematical reasoning benchmarks (zeng2025simplerlzooinvestigatingtamingzero, ; zuo2025ttrl, ; openr1, ). Additionally, we conduct an ablation study using Qwen2.5-Math-1.5B (Qwen1.5B)(yang2024qwen25mathtechnicalreportmathematical, ) to assess the effectiveness of our approach in smaller LLMs.
@@
We selected $4$ mathematical benchmarks to evaluate the models’ reasoning abilities: GSM8K (cobbe2021gsm8k, ), MATH500 (hendrycks2021measuring, ), Olympiad Bench (he2024olympiadbench, ), and College Math (tang2024mathscale, ). Each contains at least 500 data points for testing. We excluded some commonly used mathematical benchmarks that provide limited data, e,g, AIME24  [^1] https://huggingface.co/datasets/Maxwell-Jia/AIME_2024 with 30 items, as they can lead to unstable and biased evaluation outcomes. We train the base model on the GSM8K training set and then evaluate on the $4$ benchmarks.
@@
The most pertinent baselines for comparison are the base model itself and the base model trained via R1-zero. Additionally, we incorporate the latest prominent "R1-zero-Like" models with similar backbones for reference: SimpleRL-Zoo (zeng2025simplerlzooinvestigatingtamingzero, ), PRIME-Zero-7B (cui2025process, ). It is important to note that these methods are trained with different computational resources and datasets, making direct comparisons challenging. Our approach is designed to enhance diversity rather than compete directly with these methods. In fact, our method is compatible with and can be integrated into these existing approaches.
@@
[…]
@@ 5.2 Main results
We evaluate the reasoning performance using Pass@1 accuracy, as shown in Table 1. We also report the performance against training steps in Figure 4 in Appendix. In our experiment, R1-zero-Div demonstrates superior performance compared to R1-zero, achieving an average improvement of $3.5\%$. Despite being trained with limited computational resources (discussed in Appendx B), R1-zero-Div achieves comparable results to state-of-the-art methods (SimpleRL-Zoo and Eurus-2-7B-PRIME). These results suggest that promoting diversity on positive samples in training can effectively enhance the model’s reasoning capabilities. Also, following the recommendations in prior work (chandakincorrect, ; hochlehnert2025sober, ), we evaluated 8 samples per question with a temperature of $0.5$. We report Avg@8 and its standard error in the Table 2. The conclusion regarding the effectiveness of our approach remains consistent with the pass@1 metric results.
@@
We empirically demonstrate that R1-zero-Div produces more diverse solutions than other RL-finetuning baselines. Our evaluation on the GSM8K test set generates 5 responses for each of 1,319 questions, measuring diversity through three metrics: Div-Equ, and two additional metrics in prior work (kirk2023understanding, ): (1) N-gram diversity (proportion of distinct n-grams per response, capturing intra-diversity) and (2) Self-BLEU diversity (100 minus Self-BLEU score, capturing inter-diversity). All metrics range from 0 to 100, with higher values indicating greater diversity. As shown in Table 3, while RL fine-tuning methods significantly reduce diversity (compared to the base model), R1-zero-Div effectively preserves diversity. We further provide concrete examples in Appendix E showing that R1-zero-Div generates distinct solutions for the same question.
@@
Method GSM8K MATH500
Olympiad
Bench
College Math Avg
Qwen2.5-Math-7B 57.6 51.8 16.1 21.4 36.7
R1-zero 88.7 74.6 37.3 43.3 61.0
R1-zero-Div (Ours) 91.7 78.2 40.1 47.6 64.4
SimpleRL-Zoo 90.2 80.0 39.0 47.2 64.1
Eurus-2-7B-PRIME 88.0 74.4 39.4 46.6 62.1
Table 2: Avg@8 accuracy on mathematical benchmarks.
Method GSM8K MATH500
Olympiad
Bench
College Math Avg
Qwen2.5-Math-7B 53.37 (0.56) 48.10 (0.82) 15.80 (0.22) 19.36 (0.14) 34.16
R1-zero 87.77 (0.86) 72.97 (1.20) 37.26 (0.52) 42.22 (0.31) 60.06
R1-zero-Div (Ours) 90.64 (0.89) 76.92 (1.24) 39.19 (0.55) 47.49 (0.32) 63.56
SimpleRL-Zoo 89.46 (0.87) 77.15 (1.23) 39.43 (0.57) 47.19 (0.34) 63.31
Eurus-2-7B-PRIME 88.31 (0.86) 73.92 (1.18) 36.56 (0.50) 45.27 (0.30) 61.02
Table 3: Diversity of different methods on GSM8K test set.
@@
Qwen2.5-Math-7B 92.26 29.29 85.98
Eurus-2-7B-PRIME 60.86 24.08 48.20
SimpleRL-Zoo 74.89 25.41 49.32
R1-zero 75.02 27.75 56.00
zero-Div (Ours) 79.29 29.60 58.89